OpenStudy (anonymous):

Given the linear equation y = -3x – 2, find the y-coordinates of the points (0, ), (-3, ), (-1, ). Please show all of your work. Plot these points and graph the linear equation

OpenStudy (accessdenied):

Well, we know that these are the x-values for the graph. If we insert those values of x into the equation, we will find the corresponding y values. To graph, we just plot the points we find and connect them with a line. Ex.) (0,) so, x=0 y=-3x - 2 y = -3(0) - 2 y = 0 - 2 y = -2 So, our first point would be (0,-2). I think you can figure it out from here, yes? :)

OpenStudy (anonymous):

You mean the x-axis would be (0, -2)? I have grid axis and interval axis where do I write (0,-2)?

OpenStudy (accessdenied):

The point, (0,-2), like... down two units, on the y-axis.

OpenStudy (anonymous):

are you able to click this link http://resources.online.southuniversity.edu/ezgraph/media/south/Default.aspx

OpenStudy (anonymous):

You can enter any name but let me know that how to graph them on here

OpenStudy (accessdenied):

http://i44.tinypic.com/3460o6r.jpg ?

OpenStudy (anonymous):

No I have to aks my professor because they keep asking me whats the x interval and its really confusing but thansk though.

OpenStudy (accessdenied):

That's an odd question...I don't think I've ever heard of the 'x-interval' in this type of problem. Well, I'm sorry I couldn't help you. Hopefully, you can figure it out. :)

OpenStudy (anonymous):

No you did help me a lot thanks buddy,just that the stupid graph is an issue for me

OpenStudy (accessdenied):

If I had to take a guess, maybe he just means the graph's scale or the minimum x value to the maximum x-value to graph this. That would make the most sense, to me..